# Resolving power: A general approach to compare the distinguishing ability of threshold-free evaluation metrics

### v2

@@ Abstract
Selecting an evaluation metric is fundamental to model development, but uncertainty remains about when certain metrics are preferable and why. This paper introduces the concept of resolving power to describe the ability of an evaluation metric to distinguish between binary classifiers of similar quality. This ability depends on two attributes: 1. The metric’s response to improvements in classifier quality (its signal), and 2. The metric’s sampling variability (its noise). The paper defines resolving power generically as a metric’s sampling uncertainty scaled by its signal. The primary application of resolving power is to assess threshold-free evaluation metrics, such as the area under the receiver operating characteristic curve (AUROC) and the area under the precision-recall curve (AUPRC). A simulation study compares the AUROC and the AUPRC in a variety of contexts. It finds that the AUROC generally has greater resolving power, but that the AUPRC is better when searching among high-quality classifiers applied to low prevalence outcomes. The paper concludes by proposing an empirical method to estimate resolving power that can be applied to any dataset and any initial classification model.
@@ 1 Introduction
There is a large and growing collection of evaluation metrics used for binary classification models. Choosing a metric can be challenging as model evaluation serves a variety of goals. One is to provide a good description, meaning that the metric is sensitive to aspects of quality that are relevant to the user. Simple classification accuracy, for example, can be misleading if there is a large skew in the class distribution with one class occurring much more frequently than the other. Evaluation metrics are also used to select the best model from a collection of competitors [15]. This includes selection between different model classes, such as between a simple baseline model and more complex machine learning models. And it includes selection within a model class, as occurs with hyperparameter search during model tuning. Another goal of evaluation is to estimate how well a given model will perform on future, unseen cases [17].
@@
Sampling uncertainty is a neglected aspect of model evaluation within the field of machine learning [19], but it is essential to account for when data is limited [1, 5]. Metric sampling uncertainty bears upon all aspects of evaluation, though is especially important for model selection. There has been scant research that compares the sampling precision of ROC curves, PRCs, and other metrics for binary classifiers. One exception is found in Zhou [21], who used a link prediction task on a toy network model with a tunable noise parameter. For these network models Zhou finds that the area under the ROC curve (AUROC) and the area under the PRC (AUPRC) are much more discriminating than “balanced precision”, and that the AUROC is slightly more discriminating than AUPRC. [^1] Zhou finds balanced precision by choosing the decision threshold so that precision equals recall.
@@
The resolving power approach draws inspiration from Mazzanti [14], who compares how the AUROC and the AUPRC respond to improvements in classifier quality. Mazzanti’s analysis is concerned with each metric’s adequacy as a description of performance, so does not consider the issue of sampling variability. It is also limited to just one specific mechanism for model improvement. This paper provides a more general account and considers how conclusions may change under various paths to improvement. The remainder of the paper presents the resolving power methodology and then demonstrates its application for the comparison of the AUROC with the AUPRC.
@@ 2 ROC and PR curves
Our interest is in models that map cases to predicted classes. A discrete classifier is one that only outputs a class label. Applying a discrete classifier to test data produces a 2x2 confusion matrix, with rows corresponding to the predicted class and columns giving the true class. A scoring classifier outputs a number on a continuous scale, such as an estimated probability, that represents the degree to which a case belongs to a class [6]. Applying a decision threshold to a scoring classifier produces a discrete classifier. Hand [10] shows that choosing a particular threshold is equivalent to specifying the relative costs of false positives versus false negatives.
@@
The AUROC summarizes a classifier’s performance across all decision thresholds and is found by integrating the ROC curve over the $[0,1]$ range of false positive rates. Larger AUROC values are better, with the random classifier giving an $\text{AUROC}=0.5$ and the perfect classifier giving an $\text{AUROC}=1$. The AUROC, as a scalar value, is especially relevant for model tuning and selection. A disadvantage of the AUROC is that it can conceal local differences in performance. For instance, one classifier may be better for highly ranked cases while another is better for those in the intermediate or lower ranks. An important statistical property of the AUROC is that its value equals the probability that a classifier will rank a randomly chosen positive case higher than a randomly chosen negative case [9, 12]. The AUROC can be interpreted as an average sensitivity, assuming all specificity values are equally likely [10]. Several authors have demonstrated important deficiencies of the AUROC. Byrne [3] shows that it is typically not a proper scoring function. And Hand [10] argues that the AUROC is incoherent since two classifiers with the same AUROC will typically imply different relative costs of misclassification.
@@
Precision-recall (PR) graphs plot precision on the y-axis and recall on the x-axis. In PR space a random classifier corresponds to the horizontal line $y=\frac{P}{P+N}=\text{prevalence}$ where $P$ is the number of positive cases and $N$ is the number of negative cases. PR curves are sensitive to class skew while ROC curves are not. This is because inputs to the ROC curve, the true and false positive rates, only depend on the column sums of the confusion matrix. Precision depends on the row sum of true and false positives, so all else equal, it will decrease with decreasing prevalence. Insensitivity to skew has been described as both an advantage [6] and disadvantage [17] of the ROC curve.
@@
Just like the AUROC, the AUPRC reduces a scoring classifier’s performance to a single value, with larger values indicating better performance. Similar to the AUROC, the AUPRC can be interpreted as the classifier’s average precision over the $[0,1]$ range of recall values. Davis and Goadrich [4] demonstrate that there is a one-to-one correspondence between empirical ROC and PR curves since they both chart a unique mapping from confusion matrices to points in ROC or PR space. They go on to show that the AUROC and AUPRC give the same model rankings when one model’s curve “dominates” another. Informally, one curve dominates another if it lies above or equal to it across their domains. A dominating ROC curve will be northwest of the dominated curve, where its tpr is higher, its fpr is lower, or both. And a dominating PR curve will be northeast of a dominated curve, with higher precision, recall, or both across the entire domain. When there is no domination (when two curves cross) the AUROC and AUPRC can give different rankings. In cases of disagreement, the AUPRC favors classifiers with better performance in the early retrieval area, which is the region of low false positive rates in ROC space.
@@ 3 Mapping between metrics
Now we must address what we mean by “model quality”. This paper adopts a discriminative conception: Better classifiers yield greater separation in the class scores [11]. Perfect classifiers have no overlap in score distributions while totally overlapping distributions indicate a random classifier. Importantly, quality refers to the out-of-sample class score distributions, which are typically estimated with resampling methods or by using a test set.
@@
The primary challenge of this strategy is that the mapping between metrics depends on how the score distributions are ordered and we cannot know in advance how classifier improvement will change the risk scores. There are myriad ways to increase separation between distributions, each indicative of different types of improvement. This paper solves this ambiguity by fiat: It assumes that simple manipulations of score distributions are a reasonable description of model improvement. One simple approach, used below, is to add fixed increments to the positive class scores. A concern is that this mechanism may poorly describe how improvement happens in practice, but one can always explore the importance of this assumption by using alternative improvement sequences and then assessing their impact.
@@
We have identified the quality dimension as an ordered sequence of distributions, but how should we measure location on this dimension? One option is to just use the model rankings themselves, which forms an ordinal scale [18]. Another option is to measure distribution overlap directly using the Bhattacharyya coefficient. Or we can use a measure that relates overlap to model quality, the AUROC and AUPRC being two examples among many. The AUROC has several properties that make it a good choice. Since the AUROC is an area (and a probability), equal differences across the scale represent equal differences in amount. Another advantage, mentioned above, is that it is unaffected by outcome prevalence. The AUROC is also the most popular threshold-free evaluation metric for binary classifiers, making it a natural choice as the standard reference metric.
@@
For our purposes, the AUROC’s biggest advantage is that it is agnostic with respect to where changes occur in the score distributions. This fact is easiest to demonstrate with an empirical score distribution, defined as a finite set of risk scores and associated outcomes. Briefly, suppose there are $n^{+}$ positive cases, $n^{-}$ negative cases, and that all risk scores are unique. Further suppose that the classifier is not perfect, so $0.5\leq\text{AUROC}<1$, and we want to improve this by perturbing the risk scores. If we sort all cases together into a single list ranked by score, then the smallest improvements occur by finding pairs of adjacent scores that are “out-of-order”, such that the negative case has a higher score than the positive case, and re-ordering these pairs. Re-ordering a single pair will improve the AUROC by $\frac{1}{n^{+}}\times\frac{1}{n^{-}}$ regardless of where the improvement occurs. This follows from the Riemann sum method of AUROC estimation: resolving any pair of adjacent scores adds a rectangle with base $\frac{1}{n^{-}}$ and height $\frac{1}{n^{+}}$ to the area under the curve. [^3] Equivalently, on the probabilistic interpretation of the AUROC there are $n^{+}\times n^{-}$ unique ordered pairs of positive and negative scores, which forms the number of events in the sample space, and so resolving one pair increases the probability by $\frac{1}{n^{+}\times n^{-}}$. In contrast, the AUPRC will improve more for resolving out-of-order pairs that are among the highest-ranked risk scores.
@@
To summarize this section’s key points: Classifier quality is gauged by its outputs, the class score distributions. A sequence of increasingly separated class distributions forms a common quality dimension that charts the relationship between different evaluation metrics. A key caveat is that the mapping between metrics is contingent on how the score distributions are separated. Several characteristics of the AUROC make it a good choice as the reference measure of model quality. In particular, the AUROC always improves by a constant amount when resolving a pair of adjacent out-of-order risk scores. However, the viability of resolving power does not hinge on using the AUROC as the reference. Other metrics may be more appropriate for a given application.
@@ 4 Resolving power
This section illustrates the core mechanics of the approach using an idealized example while later sections move to the applications. Suppose we are interested in comparing the resolving power of the AUROC with that of the “area under the super-great curve” (the AUSGC). We construct a sampling model by specifying the class score distributions, prevalence, and sample size. Next, we create a grid of 1000 models. The first model has totally overlapping distributions for the random classifier and then we gradually shift the distributions apart so that the 1000th model has almost no overlap. Finally, we want to assess sampling models that give AUROC values of 0.7 and 0.9. We draw many replicates from these two models to estimate the sampling variability of the two metrics. Results of the analysis are summarized in Figure 2.
@@
Following the previous section’s recommendation, the AUROC on the x-axis serves as the reference scale for model quality. The left panel is then just the identity mapping. The right panel shows how the AUSGC changes relative to the AUROC, giving the relative signal of the two metrics across the quality continuum. Unit slope indicates equal signal, a slope less than 1 favors the AUROC, and a slope greater than 1 favors the AUSGC. For a given point on the curve, repeated draws from the sampling model estimates each metric’s noise distribution. The signal curves allow us to map each metric’s uncertainty interval to a model quality interval, which forms the common basis for comparison.
@@ 5 Binormal model
The binormal model, as the most commonly used in ROC analysis, serves as a good initial application of the approach. All code and simulation data used below are available on GitHub. [^4] https://github.com/colinbeam/resolving_power Now we apply the four steps of the resolving power method.
@@
Figure 4 shows relative performance, but it is also important to consider how absolute uncertainty varies across conditions. Figure 5 explores these relationships using Hanley and McNeil [12]’s formula for the approximate standard error of the AUROC. The standard error is strongly decreasing in prevalence, shown by the vertical separation between lines. The .01 condition has an especially large standard error, making relative imprecision even more costly in absolute terms. The standard error is mostly decreasing in model quality, though interestingly, slightly increases from an AUROC of 0.5 before reaching a maximum around 0.6. As an aside, we can also form “normal approximation” confidence intervals by taking plus or minus 1.96 times the standard errors from Figure 5. The normal approximation intervals are typically close to the simulation confidence intervals. For a prevalence = .01 and AUROC = .65 the simulation 95 percent confidence interval is [0.596, 0.702] while the normal approximation is [0.591, 0.709]. The approximation becomesf worse as the AUROC increases: In the prevalence = .01 and AUROC = .95 condition the simulation confidence interval is [0.929, 0.967] while the normal approximation is [0.92, 0.98]. The adequacy of this approximation bears on the utility of the alternative method for estimating resolving power, described in Appendix A.
@@
If we knew that this section’s assumptions were true—that risk scores follow a binormal distribution and model improvements come from additive shifts—then for model search it would be best to use the metric with greatest resolving power. Yet since these assumptions will never fully hold, this section’s results should be received only as general guidance to be weighed against other criteria. For instance, since the AUPRC suffers only a modest disadvantage for “good” models with moderate prevalence, it may still be preferable over the AUROC because of the greater weight it assigns to the early retrieval area.
It is uncertain how robust this section’s guidance is to deviations from the binormal model. One way to address this concern is to replace the binormal model with a domain-specific data-generating process. An example from network analysis was given above from Zhou [21]. This approach is amenable to a resolving power analysis. In fact, Zhou’s Figure 1 essentially displays signal curves by showing how evaluation metrics respond to changes in the noise parameter values that determine model quality. However, most applications will not be able to draw upon a quantitative framework to build domain-specific sampling models. An alternative strategy is to start with a dataset and a baseline model and then build the sampling model from an initial set of risk scores. This empirically-driven approach is explored in the next section.

### v3

@@ Abstract
Selecting an evaluation metric is fundamental to model development, but uncertainty remains about when certain metrics are preferable and why. This paper introduces the concept of resolving power to describe the ability of an evaluation metric to distinguish between binary classifiers of similar quality. This ability depends on two attributes: 1. The metric’s response to improvements in classifier quality (its signal), and 2. The metric’s sampling variability (its noise). The paper defines resolving power generically as a metric’s sampling uncertainty scaled by its signal. A simulation study compares the area under the receiver operating characteristic curve (AUROC) and the and the area under the precision-recall curve (AUPRC) in a variety of contexts. It finds that the AUROC generally has greater resolving power, but that the AUPRC is better when searching among high-quality classifiers applied to low prevalence outcomes. The paper also proposes an empirical method to estimate resolving power that can be applied to any dataset and any initial classification model. The AUROC is useful for developing the resolving power concept, but it has been criticized for being misleading. Newer metrics developed to address its interpretative issues can be easily incorporated into the resolving power framework. The best metrics for model search will be both interpretable and high in resolving power. Sometimes these objectives will conflict and how to address this tension remains an open question.
@@ 1 Introduction
There is a large and growing collection of evaluation metrics used for binary classification models. Choosing a metric can be challenging as model evaluation serves a variety of goals. One is interpretation, meaning that the metric is both easily understood and sensitive to aspects of quality that are relevant to the user. Simple classification accuracy, for example, is misleading for outcomes that are both low prevalence and relatively severe, such as rare but serious diseases. Evaluation metrics are also used to select the best model from a collection of competitors [15]. This includes selection between different model classes, such as between a simple baseline model and more complex machine learning models. And it includes selection within a model class, as occurs with hyperparameter search during model tuning. Another goal of evaluation is to estimate how well a given model will perform on future, unseen cases [17].
@@
Sampling uncertainty is a neglected aspect of model evaluation within the field of machine learning [19], but it is essential to account for when data is limited [1, 5]. There has been scant research that compares the sampling precision of evaluation metrics for binary classifiers. One exception is found in Zhou [21], who used a link prediction task on a toy network model with a tunable noise parameter. For these network models Zhou finds that the area under the ROC curve (AUROC) and the area under the PRC (AUPRC) are much more discriminating than “balanced precision”, and that the AUROC is slightly more discriminating than AUPRC. [^1] Zhou finds balanced precision by choosing the decision threshold so that precision equals recall.
@@
Resolving power draws inspiration from several previous works such as Saito and Rehmsmeier [17], Mazzanti [14], and hand2023notes. Each of these analyses addresses a metric’s adequacy in describing performance. In contrast, this paper’s focus on metric sampling uncertainty primarily pertains to model search and selection. The remainder of this paper presents the resolving power methodology, demonstrates its application to the AUROC and the AUPRC, and then reflects on the implications for model search.
@@ 2 ROC and PR curves
Our interest is in models that map cases to predicted classes. A discrete classifier is one that only outputs a class label. Applying a discrete classifier to test data produces a 2x2 confusion matrix (Table 1), with rows corresponding to the predicted class and columns giving the true class. A scoring classifier outputs a number on a continuous scale, such as an estimated probability, that represents the degree to which a case belongs to a class [6]. Applying a decision threshold to a scoring classifier produces a discrete classifier. Hand [10] shows that choosing a particular threshold is equivalent to specifying the relative costs of false positives versus false negatives.
@@
The AUROC summarizes a classifier’s performance across all decision thresholds and is found by integrating the ROC curve over the $[0,1]$ range of false positive rates. Larger AUROC values are better, with the random classifier giving an $\text{AUROC}=0.5$ and the perfect classifier giving an $\text{AUROC}=1$. The AUROC, as a scalar value, is especially relevant for model tuning and selection. A disadvantage of the AUROC is that it can conceal local differences in performance. For instance, one classifier may be better for highly ranked cases while another is better for those in the intermediate or lower ranks. An important statistical property of the AUROC is that its value equals the probability that a classifier will rank a randomly chosen positive case higher than a randomly chosen negative case [9, 12].
The AUROC can be interpreted as an average sensitivity, assuming all specificity values are equally likely [10]. Several authors argue that the AUROC is deficient in key respects. swamidass2010croc remark that treating all specificity values as uniformly important is not appropriate for most problems and propose the area under the concentrated ROC as a corrective. Byrne [3] shows that the AUROC is typically not a proper scoring function. The AUROC has also been criticized as incoherent since two classifiers with the same AUROC will typically imply different relative costs of misclassification [10, @hand2023notes]. Hand proposes the H-measure as a coherent alternative (Online Resource 1 applies resolving power to the H-measure). In response to Hand, ferri2011coherent argue for alternative interpretations of the AUROC that are both coherent and model independent.
@@
Precision-recall (PR) graphs plot precision on the y-axis and recall on the x-axis. In PR space a random classifier corresponds to the horizontal line $y=\frac{P}{P+N}=\text{prevalence}$ where $P$ is the number of positive cases and $N$ is the number of negative cases. PR curves are sensitive to class skew (meaning one class occurs more than the other) while ROC curves are not. This is because inputs to the ROC curve, the true and false positive rates, only depend on the column sums of the confusion matrix. Precision depends on the row sum of true and false positives, so all else equal, it will decrease with decreasing prevalence. Insensitivity to skew has been described as both an advantage [6] and disadvantage [17] of the ROC curve.
@@
Just like the AUROC, the AUPRC reduces a scoring classifier’s performance to a single value, with larger values indicating better performance. Similar to the AUROC, the AUPRC can be interpreted as the classifier’s average precision over the $[0,1]$ range of recall values. Davis and Goadrich [4] demonstrate that there is a one-to-one correspondence between empirical ROC and PR curves since they both chart a unique mapping from confusion matrices to points in ROC or PR space. They go on to show that the AUROC and AUPRC give the same model rankings when one model’s curve “dominates” another’s. Informally, one curve dominates another if it lies above or equal to it across their domains. A dominating ROC curve will be northwest of the dominated curve, where its tpr is higher, its fpr is lower, or both. And a dominating PR curve will be northeast of a dominated curve, with higher precision, recall, or both across the entire domain. When there is no domination (when two curves cross) the AUROC and AUPRC can give different rankings. In cases of disagreement, the AUPRC favors classifiers with better performance in the early retrieval area, which is the region of low false positive rates in ROC space.
@@ 3 Mapping between metrics
Now we must address what we mean by “model quality”. This paper adopts a discriminative conception: Better classifiers yield greater separation in the class scores [11]. Importantly, model quality refers to the out-of-sample class score distributions, which are typically estimated with resampling methods or by using a test set. A perfect classifier completely separates the class scores while a random classifier has identical class score distributions.333hand2023notes formally define a random classifier as one with identical cumulative distribution functions (cdfs) so that $F_{0}\equiv F_{1}$, where $F_{0}$ and $F_{1}$ are the cdfs of the negative and positive class scores, respectively. We can also use cdfs to define a perfect classifier: One where there exists a decision threshold $t^{*}$ such that $F_{0}(t^{*})=1$ and $F_{1}(t^{*})=0$. Finally, if the cdfs are not identical and they do not completely separate the class scores then they partially separate the distributions. That is, there is partial separation if $F_{0}\not\equiv F_{1}$ and there is at least one threshold $t^{\prime}$ such that $F_{0}(t^{\prime})<1$ and $F_{1}(t^{\prime})>0$. Classifiers that are neither random nor perfect are those that partially separate the score distributions.
@@
The mapping between metrics depends on how the score distributions are separated and we cannot know in advance how classifier improvement will change the risk scores. There are myriad ways to increase class separation, each indicative of different types of improvement. This paper solves this ambiguity by fiat: It assumes that simple manipulations of score distributions are a reasonable description of model improvement. A simple approach, used below, is to add fixed increments to the positive class scores. A concern is that a simple additive mechanism may poorly describe how improvement happens in practice. A more realistic, though more involved strategy is to base the mechanism on the observed early stages of model improvement (Online Resource 1 sketches this approach).
More generally, resolving power depends on two key assumptions: 1. That model improvement is a mostly homogeneous process, and 2. That we can approximate this process. The first assumption holds that improvement largely occurs in similar ways across different algorithms or hyperparameter settings. Though it does allow for random deviations due to sampling or to the intrinsic stochasticity of some algorithms. The second assumption is that we can do a decent job describing the common improvement process by finding the right sequence of class score distributions.
@@
We have identified the quality dimension as an ordered sequence of distributions, but how should we measure location on this dimension? One option is to just use the model rankings themselves, which forms an ordinal scale [18]. Another option is to measure distribution overlap directly using the Bhattacharyya coefficient. Or we can use a measure that relates overlap to model quality, the AUROC and AUPRC being two examples among many. The AUROC has several properties that make it a good choice. Since the AUROC is an area (and a probability), equal differences across the scale represent equal differences in amount. Another advantage, mentioned above, is that it is unaffected by outcome prevalence. The AUROC is also the most popular threshold-free evaluation metric for binary classifiers, making it a natural choice for the reference metric.
@@
For our purposes, the AUROC’s biggest advantage is that it is agnostic with respect to where changes occur in the score distributions. This fact is easiest to demonstrate with an empirical score distribution, defined as a finite set of risk scores and associated outcomes. Briefly, suppose there are $n^{+}$ positive cases, $n^{-}$ negative cases, and that all risk scores are unique. Further suppose that the classifier is not perfect, so $0.5\leq\text{AUROC}<1$, and we want to improve this by perturbing the risk scores. If we sort all cases together into a single list ranked by score, then the smallest improvements occur by finding pairs of adjacent scores that are “out-of-order”, such that the negative case has a higher score than the positive case, and re-ordering these pairs. Re-ordering a single pair will improve the AUROC by $\frac{1}{n^{+}}\times\frac{1}{n^{-}}$ regardless of where the improvement occurs. This follows from the probabilistic interpretation of the AUROC: There are $n^{+}\times n^{-}$ unique ordered pairs of positive and negative scores, which forms the number of events in the sample space. So resolving one out-of-order pair increases the probability by $\frac{1}{n^{+}\times n^{-}}$. In contrast, the AUPRC will improve more for resolving out-of-order pairs that are among the highest-ranked risk scores.
@@
To summarize this section’s key points: Classifier quality is gauged by its outputs, the class score distributions. A sequence of increasingly separated class distributions forms a common quality dimension that charts the relationship between different evaluation metrics. A key caveat is that the mapping between metrics is contingent on how the score distributions are separated. Several characteristics of the AUROC make it a good choice as the reference measure of model quality. [^4] Note that these advantages pertain narrowly to the AUROC’s role as reference measure and do not counter to the criticisms of the AUROC referenced in the previous section. In particular, the AUROC always improves by a constant amount when resolving a pair of adjacent out-of-order risk scores. Moreover, the viability of resolving power does not hinge on using the AUROC as the reference. Instead, it is the sequence of class score distributions that are fundamental. Other metrics can be substituted so long as they form a mapping from the class score distributions to an interval scale.
@@ 4 Resolving power
This section illustrates the core mechanics of the approach using an idealized example while later sections move to the applications. Suppose we are interested in comparing the resolving power of the AUROC with that of the fictional “area under the super-great curve” (the AUSGC). We construct a sampling model by specifying the class score distributions, prevalence, and sample size. Next, we create a grid of 1000 models. The first model has identical class score distributions for the random classifier and then we gradually shift the distributions apart so that the 1000th model has almost no overlap. Finally, we want to assess sampling models that give AUROC values of 0.7 and 0.9. We draw many replicates from these two models to estimate the sampling variability of the two metrics. Figure 2 summarizes the analysis.
@@
Following the previous section’s recommendation, the AUROC on the x-axis serves as the reference scale for model quality. The left panel is then just the identity mapping. The right panel shows how the AUSGC changes relative to the AUROC, giving the relative signal of the two metrics across the quality continuum. Unit slope indicates equal signal, a slope less than 1 favors the AUROC, and a slope greater than 1 favors the AUSGC. For a given point on the curve, repeated draws from the sampling model estimate each metric’s noise distribution. The signal curves allow us to map each metric’s uncertainty interval to a model quality interval, which forms the common basis for comparison.
@@ 5 Binormal model
The binormal model, as the most commonly used in ROC analysis, serves as a good initial application of the approach. All code and data used below are available on GitHub. [^5] https://github.com/colinbeam/resolving_power Now we apply the four steps of the resolving power method.
@@
Figure 4 shows relative performance, but it is also important to consider how absolute uncertainty varies across conditions. Figure 5 explores these relationships using Hanley and McNeil [12]’s formula for the approximate standard error of the AUROC. The standard error is strongly decreasing in prevalence, shown by the vertical separation between lines. The .01 condition has an especially large standard error, making relative imprecision even more costly in absolute terms. The standard error is mostly decreasing in model quality, though interestingly, slightly increases from an AUROC of 0.5 before reaching a maximum around 0.6. As an aside, we can also form “normal approximation” confidence intervals by taking plus or minus 1.96 times the standard errors from Figure 5. The normal approximation intervals are typically close to the simulation confidence intervals. For a prevalence = .01 and AUROC = .65 the simulation 95 percent confidence interval is [0.596, 0.702] while the normal approximation is [0.591, 0.709]. The approximation becomes worse as the AUROC increases: In the prevalence = .01 and AUROC = .95 condition the simulation confidence interval is [0.929, 0.967] while the normal approximation is [0.92, 0.98]. The adequacy of this approximation bears on the utility of the alternative method for estimating resolving power, described in Appendix A.
@@
Now suppose the AUPRC is what we really want to maximize. That is, the AUPRC best captures our intuitions about what makes a good model. But also suppose that we know our resolving power assumptions are true: That risk scores follow a binormal distribution and model improvements come from additive shifts. Then what this section’s results tell us is that we should use the AUROC for search even if our goal is to find the best AUPRC model (at least in most contexts). Under this hypothetical scenario we don’t need to balance the AUROC’s greater resolving power against our preference for the AUPRC because we know precisely when the AUROC is a better guide. But once we select the final model we should still describe its performance with the AUPRC since that is what we really care about.
Moving from the hypothetical to reality, we know that this section’s assumptions will never fully hold. This means that the results should be taken as only general guidance to be weighed against other criteria. For instance, the AUPRC suffers only a modest disadvantage for “good” models with moderate prevalence. Once we factor the influence of assumptions violations, which we could test with sensitivity analyses that explore other paths towards class score separation, the disadvantage may disappear.
It is uncertain how robust this section’s guidance is to deviations from the binormal model. One way to address this concern is to replace the binormal with a domain-specific data-generating process. However, most applications will not have the requisite quantitative framework to accomplish this. An alternative strategy is to start with a dataset and a baseline model and then build the sampling model from an initial set of risk scores. This empirically-driven approach is explored in the next section.